# Global AI Governance Overview: Understanding Regulatory Requirements Across Global Jurisdictions

### v1

@@ 2 Research Methodology
The research methodology employed in this study consists of a systematic documentary analysis approach. The documentation that was examined includes AICM Framework [csa2025aicm], AI Privacy Risks & Mitigations in LLM [edpb2025privacy], General Data Protection Regulation (hereinafter GDPR) [gdpr2016], AI Code of Practice [eu2024codeofpractice], EU AI Act [euaiact2024], US AI Action Plan [whitehouse2025aiplan], NIST AI RMF 1.0 [nist2023airmf], OWASP Top 10 LLM Vulnerabilities [owasp2025top10], ISO/IEC 42005:2025 [iso42005:2025], Digital Services Act (hereinafter DSA) [dsa2022] and AI TRiSM [ibm2024trism], Directive 2001/29/EC [eucopyright2001consolidated], Council Regulation (EC) No 1383/2003 [eucustoms2003], Directive (EU) 2019/790 [eucopyright2019], California Senate Bill 1001 [california_sb1001_2018], California Senate Bill 942 [california_sb942_2024], Utah Senate Bill 149 [utah_sb149_2024], Arkansas House Bill 1958 [arkansas_hb1958_2025], Arkansas Act 927 [arkansas_act927_2025], Maryland House Bill 956 [maryland_hb956_2025], Montana Senate Bill 212 [montana_sb212_2025], Illinois House Bill 3773 [illinois_hb3773_2025], West Virginia House Bill 3187 [westvirginia_hb3187_2025], China’s Cybersecurity Law [creemers2018translation], China’s Personal Information Protection Law [PIPL2021], Administrative Provisions on Algorithmic Recommendation Services [china2022algorithmic], Provisions on the Administration of Deep Synthesis Internet Information Services [china2022deepfakes], Interim Measures for the Management of Generative AI Services [china2023generativeai], Japan’s Social Principles of Human-Centric AI [cabinet2019humancentricAI], AI Governance Guidelines for Business [meti2024aiguidelinesV1], AI Governance Framework [grjapan2024aiJapan], South Korea’s Basic Act on AI [korea2024AIBasicAct], Personal Information Protection Act (PIPA) [koreaPIPA], Network Act [koreaNetworkAct], and Product Liability Act [koreaProductLiabilityAct].
@@ 3.1 Terminology
General-Purpose AI Models (hereinafter GPAI) are defined by their significant generality and ability to competently execute diverse tasks [euaiact2024, Article 3 (63)]. These models, often trained with substantial data volumes using self-supervision techniques, are designed for integration across multiple downstream systems and applications.
Signatories refer to major AI companies that provide large-scale AI models or AI-powered services and that have committed to voluntary agreements with regulatory bodies, for example entities such as OpenAI and Google [eu2024codeofpractice].
Downstream providers are organizations that integrate general-purpose AI models into their own AI systems or applications [eu2024codeofpractice]. This includes entities that: (1) fine-tune or modify GPAI models, (2) build applications that incorporate GPAI model functionality beyond simple API access, (3) create AI systems where GPAI models perform core functions, or (4) develop products that combine GPAI models with additional AI components. Simple API wrappers or user interfaces that only facilitate access to unmodified GPAI models without additional AI functionality are not considered downstream providers under this definition [euaiact2024, Article 3(68)].
The AI Office is the European Commission’s specialized function responsible for implementing, monitoring, and supervising AI systems and GPAI models, as well as overseeing AI governance [euaiact2024, Article 3(47)]. All references to the AI Office in the regulation are understood as references to the European Commission itself.
Deployer is an entity that uses AI systems within their operations, excluding personal use [euaiact2024, Article 3(4)]. This includes companies, government agencies, and organizations that implement AI systems in their work.
Provider is an organization that develops AI systems or GPAI models and distributes them under their own brand. [euaiact2024, Article 3(3)]. This includes both direct development and commissioning others to develop AI on their behalf.
Distributor is an entity that supplies AI systems to the EU market without developing or importing them [euaiact2024, Article 3(7)]. Examples include software resellers, technology retailers, or platforms that sell third-party AI solutions.
Systemic risk means a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact on the European Union market due to their reach, or due to actual or reasonably foreseeable negative effects on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain [euaiact2024, Article 3(65)].
Model Documentation Form is a form used to describe all the details about how the GPAI works, including all the technical details such as training process, computational resources, and energy consumption [eu2024modeldocform].
@@ 3.2.1 Regulation Analysis
  • EU AI Act is a regulation aimed at establishing a framework for the development and use of artificial intelligence on the territory of the European Union [euaiact2024]. The implementation and compliance with EU AI Act is controlled by the AI Office mentioned before.
  • GDPR is an EU law that sets out rules for how organizations collect, use and protect personal data of individuals within the EU [gdpr2016]. The connection between GDPR and the EU AI Act lies in the fact that GDPR affects the data on which AI systems can be trained, protecting intellectual property and personal identifiable information of individuals.
  • AI Code of Practice is a voluntary set of guidelines designed to help providers of GPAI models to comply with the EU AI Act [eu2024codeofpractice]. It plays a crucial role in the governance of AI systems.
  • DSA is a European Union law focused on regulating online platforms and digital services [dsa2022]. Platforms that use AI for content moderation or advertising must ensure that their AI systems meet the standards of both the DSA and EU AI Act.
@@ 3.2.2 EU AI Act: Covered Entities
Article 53 of EU AI Act states that open-source AI models (specifically GPAI) released under free and open-source licenses do receive exemptions from some AI Act obligations (e.g providing documentation to downstream providers, or upon request, to the AI Office or national authorities), but these exemptions do not apply if the model is classified as having systemic risk [euaiact2024, Article 53]. In such case, they still must comply with training data summary and copyright policy requirements.
@@ 3.3 EU AI Code of Practice
The EU AI Code of Practice [eu2024codeofpractice] document is split into three main categories, which are Transparency, Copyright, Safety and Security. This section undertakes a systematic review of the principal implementation guidelines, examining the specific entities and actors to whom these provisions are addressed. While the EU AI Act establishes binding regulatory obligations, the Code of Practice provides voluntary guidance to demonstrate compliance with these requirements.
@@ 3.3.1 Transparency Chapter
The Transparency chapter proposes three measures that serve as guidance for demonstrating compliance with transparency obligations under the AI Act [euaiact2024, Article 53, Article 55], specifically addressing Articles 53 and 55 requirements [eu2024codeofpractice, Transparency Chapter, Objectives].
Measure 1.1 [eu2024codeofpractice, Transparency Chapter, Measure 1.1]
@@
Measure 1.2 [eu2024codeofpractice, Transparency Chapter, Measure 1.2]
@@
Measure 1.3 [eu2024codeofpractice, Transparency Chapter, Measure 1.3]
Signatories must ensure that Model Documentation is controlled for integrity, quality and is protected from unwanted alterations. However, it is important to note that this chapter does not directly affect Downstream Providers, since the Model Documentation Form [eu2024modeldocform] is filled in by Signatories.
@@ 3.3.2 Copyright Chapter
Copyright chapter presents that Signatories should have a copyright policy and controls that correspond to EU Copyright law, which is described in several EU Directives [eucopyright2019] such as Directive 2001/29/EC on the harmonisation of copyright and related rights in the information society [eucopyright2001consolidated] and Directive (EU) 2019/790 on copyright and related rights in the Digital Single Market [eucopyright2019]. Signatories are responsible for verifying that all their measures of copyright are compliant in all the EU member countries.
Measure 1.1 [eu2024codeofpractice, Copyright Chapter, Measure 1.1]
@@
Measure 1.2 [eu2024codeofpractice, Copyright Chapter, Measure 1.2]
If the technological denial or other preventing from scraping techniques or signs installed on the web, the Signatories must not try to scrape that information. Also, there is a list with links to resources which must not be touched (for example piracy websites, sites with massive copyright infringement, or the ones that undergo the court process). This list is necessary to help AI companies avoid training their models on content from websites that have been officially recognized by courts or public authorities as systematically violating copyright laws, thereby reducing the risk of incorporating infringing material into their training datasets and subsequently generating copyright-violating outputs. However, this dynamic list has not yet been made publicly available by the European Commission, though it is planned to be published on an EU website [europeancommission2024].
The Signatories must make this information available (on their website for example) and state how their web crawling system works in the technical documentation (Model Documentation Form). The requirements for this form can be found in the Annex IV 2(d) of the EU AI Act [euaiact2024, Annex IV].
Measure 1.3 [eu2024codeofpractice, Copyright Chapter, Measure 1.3]
@@
Measure 1.4 [eu2024codeofpractice, Copyright Chapter, Measure 1.4]
Signatories must prevent their models from generating infringing outputs when the downstream provider implements their model. It should be stated in copyright policy, terms, and conditions. Or if it is a model released under open-source license then the prohibition of copyright infringing uses must be stated in the accompanying documentation. The infringing rights are defined in Council Regulation (EC) No 1383/2003 of 22 July 2003 [eucustoms2003]. It directly specifies that these are considered infringement:
@@
Measure 1.5 [eu2024codeofpractice, Copyright Chapter, Measure 1.5]
@@
The Code of Practice does not specify what exactly must be done with the data (deleted or else). Then , the Directive2001/29/EC must be considered, specifically Article 8 [eucopyright2001consolidated, Article 8], which states:
@@
According to Article 8 the material must be seizured. Also, Article 17 of GDPR [gdpr2016, Article 7] states the “right to be forgotten”, which brings a concern with model unlearning methodology.
The commitment of data removal does not affect the measures, remedies and sanctions available to enforce copyright and related rights under Union and national law [eu2024codeofpractice].
@@ 3.3.3 Safety and Security Chapter
Measure 1.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 1.1]
@@
Measure 1.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 1.2]
@@
Measure 1.3 [eu2024codeofpractice, Safety and Security Chapter, Measure 1.3]
@@
Measure 1.4 [eu2024codeofpractice, Safety and Security Chapter, Measure 1.4]
@@
Measure 2.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 2.1]
@@
Measure 2.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 2.2]
@@
Measure 3.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 3.1]
@@
Measure 3.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 3.2]
@@
Measure 3.3 [eu2024codeofpractice, Safety and Security Chapter, Measure 3.3]
@@
Measure 3.4 [eu2024codeofpractice, Safety and Security Chapter, Measure 3.4]
@@
Measure 3.5 [eu2024codeofpractice, Safety and Security Chapter, Measure 3.5]
@@
Measure 4.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 4.1]
@@
Measure 4.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 4.2]
@@
Measure 5.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 5.1]
@@
Measure 6.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 6.1]
@@
Measure 6.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 6.2]
@@
Measure 7.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 7.1]
@@
Measure 7.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 7.2]
@@
Measure 7.3 [eu2024codeofpractice, Safety and Security Chapter, Measure 7.3]
@@
Measure 7.4 [eu2024codeofpractice, Safety and Security Chapter, Measure 7.4]
@@
Measure 7.5 [eu2024codeofpractice, Safety and Security Chapter, Measure 7.5]
@@
Measure 7.6 [eu2024codeofpractice, Safety and Security Chapter, Measure 7.6]
@@
Measure 7.7 [eu2024codeofpractice, Safety and Security Chapter, Measure 7.7]
@@
Measure 8.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 8.1]
@@
Measure 8.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 8.2]
@@
Measure 8.3 [eu2024codeofpractice, Safety and Security Chapter, Measure 8.3]
@@
Measure 9.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 9.1]
@@
Measure 9.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 9.2]
@@
Measure 9.3 [eu2024codeofpractice, Safety and Security Chapter, Measure 9.3]
@@
Measure 9.4 [eu2024codeofpractice, Safety and Security Chapter, Measure 9.4]
@@
Measure 10.1 [eu2024codeofpractice, Safety and Security Chapter, Measure 10.1]
@@
Measure 10.2 [eu2024codeofpractice, Safety and Security Chapter, Measure 10.2]

### v2

@@ 2 Research Methodology
The research methodology employed in this study consists of a systematic documentary analysis approach. The documentation that was examined includes AICM Framework [6], AI Privacy Risks & Mitigations in LLM [13], General Data Protection Regulation (hereinafter GDPR) [15], AI Code of Practice [10], EU AI Act [18], US AI Action Plan [41], NIST AI RMF 1.0 [36], OWASP Top 10 LLM Vulnerabilities [37], ISO/IEC 42005:2025 [28], Digital Services Act (hereinafter DSA) [17] and AI TRiSM [23], Directive 2001/29/EC [14], Council Regulation (EC) No 1383/2003 [7], Directive (EU) 2019/790 [16], California Senate Bill 1001 [4], California Senate Bill 942 [5], Utah Senate Bill 149 [42], Arkansas House Bill 1958 [2], Arkansas Act 927 [1], Maryland House Bill 956 [32], Montana Senate Bill 212 [35], Illinois House Bill 3773 [24], West Virginia House Bill 3187 [43], China’s Cybersecurity Law [8], China’s Personal Information Protection Law [9], Administrative Provisions on Algorithmic Recommendation Services [19], Provisions on the Administration of Deep Synthesis Internet Information Services [20], Interim Measures for the Management of Generative AI Services [21], Japan’s Social Principles of Human-Centric AI [3], AI Governance Guidelines for Business [33], AI Governance Framework [22], South Korea’s Basic Act on AI [34], Personal Information Protection Act (PIPA) [30], Network Act [29], and Product Liability Act [31].
@@ 3.1 Terminology
General-Purpose AI Models (hereinafter GPAI) are defined by their significant generality and ability to competently execute diverse tasks [18, Article 3 (63)]. These models, often trained with substantial data volumes using self-supervision techniques, are designed for integration across multiple downstream systems and applications.
Signatories refer to major AI companies that provide large-scale AI models or AI-powered services and that have committed to voluntary agreements with regulatory bodies, for example entities such as OpenAI and Google [10].
Downstream providers are organizations that integrate general-purpose AI models into their own AI systems or applications [10]. This includes entities that: (1) fine-tune or modify GPAI models, (2) build applications that incorporate GPAI model functionality beyond simple API access, (3) create AI systems where GPAI models perform core functions, or (4) develop products that combine GPAI models with additional AI components. Simple API wrappers or user interfaces that only facilitate access to unmodified GPAI models without additional AI functionality are not considered downstream providers under this definition [18, Article 3(68)].
The AI Office is the European Commission’s specialized function responsible for implementing, monitoring, and supervising AI systems and GPAI models, as well as overseeing AI governance [18, Article 3(47)]. All references to the AI Office in the regulation are understood as references to the European Commission itself.
Deployer is an entity that uses AI systems within their operations, excluding personal use [18, Article 3(4)]. This includes companies, government agencies, and organizations that implement AI systems in their work.
Provider is an organization that develops AI systems or GPAI models and distributes them under their own brand. [18, Article 3(3)]. This includes both direct development and commissioning others to develop AI on their behalf.
Distributor is an entity that supplies AI systems to the EU market without developing or importing them [18, Article 3(7)]. Examples include software resellers, technology retailers, or platforms that sell third-party AI solutions.
Systemic risk means a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact on the European Union market due to their reach, or due to actual or reasonably foreseeable negative effects on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain [18, Article 3(65)].
Model Documentation Form is a form used to describe all the details about how the GPAI works, including all the technical details such as training process, computational resources, and energy consumption [11].
@@ 3.2.1 Regulation Analysis
  • EU AI Act is a regulation aimed at establishing a framework for the development and use of artificial intelligence on the territory of the European Union [18]. The implementation and compliance with EU AI Act is controlled by the AI Office mentioned before.
  • GDPR is an EU law that sets out rules for how organizations collect, use and protect personal data of individuals within the EU [15]. The connection between GDPR and the EU AI Act lies in the fact that GDPR affects the data on which AI systems can be trained, protecting intellectual property and personal identifiable information of individuals.
  • AI Code of Practice is a voluntary set of guidelines designed to help providers of GPAI models to comply with the EU AI Act [10]. It plays a crucial role in the governance of AI systems.
  • DSA is a European Union law focused on regulating online platforms and digital services [17]. Platforms that use AI for content moderation or advertising must ensure that their AI systems meet the standards of both the DSA and EU AI Act.
@@ 3.2.2 EU AI Act: Covered Entities
Article 53 of EU AI Act states that open-source AI models (specifically GPAI) released under free and open-source licenses do receive exemptions from some AI Act obligations (e.g providing documentation to downstream providers, or upon request, to the AI Office or national authorities), but these exemptions do not apply if the model is classified as having systemic risk [18, Article 53]. In such case, they still must comply with training data summary and copyright policy requirements.
@@ 3.3 EU AI Code of Practice
The EU AI Code of Practice [10] document is split into three main categories, which are Transparency, Copyright, Safety and Security. This section undertakes a systematic review of the principal implementation guidelines, examining the specific entities and actors to whom these provisions are addressed. While the EU AI Act establishes binding regulatory obligations, the Code of Practice provides voluntary guidance to demonstrate compliance with these requirements.
@@ 3.3.1 Transparency Chapter
The Transparency chapter proposes three measures that serve as guidance for demonstrating compliance with transparency obligations under the AI Act [18, Article 53, Article 55], specifically addressing Articles 53 and 55 requirements [10, Transparency Chapter, Objectives].
Measure 1.1 [10, Transparency Chapter, Measure 1.1]
@@
Measure 1.2 [10, Transparency Chapter, Measure 1.2]
@@
Measure 1.3 [10, Transparency Chapter, Measure 1.3]
Signatories must ensure that Model Documentation is controlled for integrity, quality and is protected from unwanted alterations. However, it is important to note that this chapter does not directly affect Downstream Providers, since the Model Documentation Form [11] is filled in by Signatories.
@@ 3.3.2 Copyright Chapter
Copyright chapter presents that Signatories should have a copyright policy and controls that correspond to EU Copyright law, which is described in several EU Directives [16] such as Directive 2001/29/EC on the harmonisation of copyright and related rights in the information society [14] and Directive (EU) 2019/790 on copyright and related rights in the Digital Single Market [16]. Signatories are responsible for verifying that all their measures of copyright are compliant in all the EU member countries.
Measure 1.1 [10, Copyright Chapter, Measure 1.1]
@@
Measure 1.2 [10, Copyright Chapter, Measure 1.2]
If the technological denial or other preventing from scraping techniques or signs installed on the web, the Signatories must not try to scrape that information. Also, there is a list with links to resources which must not be touched (for example piracy websites, sites with massive copyright infringement, or the ones that undergo the court process). This list is necessary to help AI companies avoid training their models on content from websites that have been officially recognized by courts or public authorities as systematically violating copyright laws, thereby reducing the risk of incorporating infringing material into their training datasets and subsequently generating copyright-violating outputs. However, this dynamic list has not yet been made publicly available by the European Commission, though it is planned to be published on an EU website [12].
The Signatories must make this information available (on their website for example) and state how their web crawling system works in the technical documentation (Model Documentation Form). The requirements for this form can be found in the Annex IV 2(d) of the EU AI Act [18, Annex IV].
Measure 1.3 [10, Copyright Chapter, Measure 1.3]
@@
Measure 1.4 [10, Copyright Chapter, Measure 1.4]
Signatories must prevent their models from generating infringing outputs when the downstream provider implements their model. It should be stated in copyright policy, terms, and conditions. Or if it is a model released under open-source license then the prohibition of copyright infringing uses must be stated in the accompanying documentation. The infringing rights are defined in Council Regulation (EC) No 1383/2003 of 22 July 2003 [7]. It directly specifies that these are considered infringement:
@@
Measure 1.5 [10, Copyright Chapter, Measure 1.5]
@@
The Code of Practice does not specify what exactly must be done with the data (deleted or else). Then , the Directive2001/29/EC must be considered, specifically Article 8 [14, Article 8], which states:
@@
According to Article 8 the material must be seizured. Also, Article 17 of GDPR [15, Article 7] states the “right to be forgotten”, which brings a concern with model unlearning methodology.
The commitment of data removal does not affect the measures, remedies and sanctions available to enforce copyright and related rights under Union and national law [10].
@@ 3.3.3 Safety and Security Chapter
Measure 1.1 [10, Safety and Security Chapter, Measure 1.1]
@@
Measure 1.2 [10, Safety and Security Chapter, Measure 1.2]
@@
Measure 1.3 [10, Safety and Security Chapter, Measure 1.3]
@@
Measure 1.4 [10, Safety and Security Chapter, Measure 1.4]
@@
Measure 2.1 [10, Safety and Security Chapter, Measure 2.1]
@@
Measure 2.2 [10, Safety and Security Chapter, Measure 2.2]
@@
Measure 3.1 [10, Safety and Security Chapter, Measure 3.1]
@@
Measure 3.2 [10, Safety and Security Chapter, Measure 3.2]
@@
Measure 3.3 [10, Safety and Security Chapter, Measure 3.3]
@@
Measure 3.4 [10, Safety and Security Chapter, Measure 3.4]
@@
Measure 3.5 [10, Safety and Security Chapter, Measure 3.5]
@@
Measure 4.1 [10, Safety and Security Chapter, Measure 4.1]
@@
Measure 4.2 [10, Safety and Security Chapter, Measure 4.2]
@@
Measure 5.1 [10, Safety and Security Chapter, Measure 5.1]
@@
Measure 6.1 [10, Safety and Security Chapter, Measure 6.1]
@@
Measure 6.2 [10, Safety and Security Chapter, Measure 6.2]
@@
Measure 7.1 [10, Safety and Security Chapter, Measure 7.1]
@@
Measure 7.2 [10, Safety and Security Chapter, Measure 7.2]
@@
Measure 7.3 [10, Safety and Security Chapter, Measure 7.3]
@@
Measure 7.4 [10, Safety and Security Chapter, Measure 7.4]
@@
Measure 7.5 [10, Safety and Security Chapter, Measure 7.5]
@@
Measure 7.6 [10, Safety and Security Chapter, Measure 7.6]
@@
Measure 7.7 [10, Safety and Security Chapter, Measure 7.7]
@@
Measure 8.1 [10, Safety and Security Chapter, Measure 8.1]
@@
Measure 8.2 [10, Safety and Security Chapter, Measure 8.2]
@@
Measure 8.3 [10, Safety and Security Chapter, Measure 8.3]
@@
Measure 9.1 [10, Safety and Security Chapter, Measure 9.1]
@@
Measure 9.2 [10, Safety and Security Chapter, Measure 9.2]
@@
Measure 9.3 [10, Safety and Security Chapter, Measure 9.3]
@@
Measure 9.4 [10, Safety and Security Chapter, Measure 9.4]
@@
Measure 10.1 [10, Safety and Security Chapter, Measure 10.1]
@@
Measure 10.2 [10, Safety and Security Chapter, Measure 10.2]